# Spectropolarimeter on a 2–4 m class telescope and proposed science cases

### v1

@@ 4 Calibration, Data reduction and polarization estimation methods
Data reduction will be done using standard data reduction routines available to the community. Calibration of wide-field polarimetry instruments is a challenging problem owing to the combination of the following two effects: (a) for wide-field instruments, there is large change in the polarimetric response function of the instrument across the FoV, requiring wide-field calibration sources that can fill the entire FoV, and (b) currently, prior to the work by the authors, there have been no established wide-field calibration sources. A more detailed discussion on this subject can be found in the polarimetric modelling and calibration paper for WALOP instruments by Maharana et al (ref. 21). They have developed a high accuracy polarimetry calibration method for wide-field polarimeters, in particular for WALOP-South instrument.

### v2

@@ 4 Calibration, Data reduction and polarization estimation methods
Data reduction will be done using standard data reduction routines available to the community. Calibration of wide-field polarimetry instruments is a challenging problem owing to the combination of the following two effects: (a) for wide-field instruments, there is large change in the polarimetric response function of the instrument across the FoV, requiring wide-field calibration sources that can fill the entire FoV, and (b) currently, prior to the work by the authors22,23, there have been no established wide-field calibration sources. A more detailed discussion on this subject can be found in the polarimetric modelling and calibration paper for WALOP instruments by Maharana et al (ref. 21). They have developed a high accuracy polarimetry calibration method for wide-field polarimeters, in particular for WALOP-South instrument.